one mole of an ideal gas is held at a constant volume of 1 lt. find the change in pressure if the temperature increases by 50 c
Answer #1

Answer:

1 litre of volume stands for 0.001m3

For an ideal gas, PV=nRT stands at all times.

So, for state 1, P1V1=nRT1 ........1)

and for state 2, P2V2=nRT2 ........2)

Subtracting 1) from 2) gives us

P2V2 - P1V1= nRT2 - nRT1

Since, volume is constant

V(P2-P1) = nR(T2-T1)

So, (P2-P1) = nR(T2-T1)/V = 1*8.314*50/0.001 = 415700 Pa